# Partial Inverse Design of High-Performance Concrete Using Cooperative Neural Networks for Constraint-Aware Mix Generation

### v1

@@ Abstract
High-performance concrete offers exceptional strength and durability but requires complex mix designs involving many interdependent variables and practical constraints. While data-driven methods have advanced predictive modeling for forward design, inverse design, which focuses on determining mix compositions that achieve target performance, remains limited, particularly in design situations where some mix variables are fixed by constraints and only the remaining variables must be determined. This study proposes a cooperative neural network framework for the partial inverse design of high-performance concrete. The framework combines two coupled neural network models, an imputation model that infers the undetermined variables and a surrogate model that predicts compressive strength. Through cooperative learning, the model generates valid and performance-consistent mix designs in a single forward pass while accommodating different constraint combinations without retraining. Its performance is compared with both probabilistic and generative approaches, including Bayesian inference based on a Gaussian process surrogate and autoencoder-based models. Evaluated on a benchmark dataset, the proposed model achieves stable and higher R-squared values of 0.87-0.92 and reduces mean squared error by an average of 50 percent compared with autoencoder baselines and by an average of 70 percent compared with Bayesian inference. The results demonstrate that the cooperative neural network provides an accurate, robust, and computationally efficient foundation for constraint-aware, data-driven mix proportioning in concrete engineering.

### v2

@@ Abstract
High-performance concrete requires complex mix design decisions involving interdependent variables and practical constraints. While data-driven methods have improved predictive modeling for forward design in concrete engineering, inverse design remains limited, especially when some variables are fixed and only the remaining ones must be inferred. This study proposes a cooperative neural network framework for the partial inverse design of high-performance concrete. The framework integrates an imputation model with a surrogate strength predictor and learns through cooperative training. Once trained, it generates valid and performance-consistent mix designs in a single forward pass without retraining for different constraint scenarios. Compared with baseline models, including autoencoder models and Bayesian inference with Gaussian process surrogates, the proposed method achieves R-squared values of 0.87 to 0.92 and substantially reduces mean squared error by approximately 50% and 70%, respectively. The results show that the framework provides an accurate and computationally efficient foundation for constraint-aware, data-driven mix proportioning.
@@ 4.3 Perspective on Model Application
Category Model Design variables Estimation strength (diff)
Cement BFS PFA Water SP CA FA Age
User input 212.50 155.70 14.30 880.40 28.00 55.50
Modelling results Bayesian-GP 372.54 212.50 89.56 155.70 14.30 943.51 880.40 28.00 69.81 (+14.31)
Standalone DAE 381.83 212.50 20.12 155.70 14.30 957.71 880.40 28.00 70.67 (+15.18)
Standalone DVAE 377.34 212.50 35.61 155.70 14.30 915.17 880.40 28.00 71.67 (+16.18)
Standalone DWAE 429.14 212.50 2.97 155.70 14.30 828.02 880.40 28.00 71.94 (+16.44)
Cooperative DAE 294.17 212.50 44.94 155.70 14.30 971.42 880.40 28.00 61.72 (+6.23)
Cooperative DVAE 265.80 212.50 65.40 155.70 14.30 961.96 880.40 28.00 59.72 (+4.22)
Cooperative DWAE 283.38 212.50 88.24 155.70 14.30 951.90 880.40 28.00 60.91 (+5.41)